OpenStudy (anonymous):

What quantity of 70 per cent acid solution must be mixed with a 25 solution to produce 540 mL of a 50 per cent solution?

OpenStudy (amistre64):

you seem to have alot of these, how are you ever to go thru them efficiently?

OpenStudy (amistre64):

i got a formula worked out for these

OpenStudy (anonymous):

im hoping someone will tell me a formula to help me out, most ppl on here usually do that or show me how to get the answer

OpenStudy (amistre64):

s + t = 540 ; *-.25 s(.7) + t(.25) = 540(.5) -s(.25) - t(.25)= 540(-.25) s(.7) + t(.25) = 540(.5) -------------------------- s(.7-.25) = 540(.5-.25) s = 540(.5-.25)/(.7-.25)

OpenStudy (anonymous):

we haven't went over this in class yet, we start word problems next week, these are practice problems given to see how much we know

OpenStudy (amistre64):

that gives me 300 of 70%

OpenStudy (anonymous):

why is it -.25

OpenStudy (amistre64):

to get rid of the t(.25) we need to multiply the top by -.25

OpenStudy (amistre64):

you see how it cancels out?

OpenStudy (amistre64):

i read a book from my library called something like: how to solve word problems. becasue tey used to be and sometimes still are my weak point

OpenStudy (anonymous):

who is the author?

OpenStudy (amistre64):

cant recall, some old dead guy i think

OpenStudy (anonymous):

lol, yeah i've never been good at word problems, i'm even worse at percentages and when you put them together it's a little hard for me to comprehend until the light bulb comes on and then i get it

OpenStudy (anonymous):

but math has always been that way for me, i just sit in class and try to follow until the light bulb comes on

OpenStudy (amistre64):

i was like that in stats class, memorizing formulas aint for me; and about 2 weeks in i started to understand it better

OpenStudy (anonymous):

can you help me with my other problem about fixed cost and unit cost.... im not sure what fixed cost means....